solve
Vy+5 -1=3
the v is suppose to represent a square root (ex:the square root of y+5)

Answer by nyc_function(2741)

sqrt{y+5} -1 = 3

Isolate the radical as step one.

sqrt{y + 5} = 3 + 1

sqrt{y + 5} = 4

We now square both sides of the equation. We must square in order to remove the radical from the equation.

[sqrt{y + 5}]^2 = (4)^2

y + 5 = 16

y = 16 - 5

y = 11

We now plug y = 11 into the given question.
The goal is to get the same answer on both sides of the equation.

sqrt{y+5} -1 = 3
Let y = 11
sqrt{11+5} -1 = 3

sqrt{16} - 1 = 3

4 - 1 = 3
3 = 3...It checks!!

The right answer is y = 11.

Done!